OpenStudy (anonymous):

Find the sum: -16 + (-5) = A. -21 B. -11 C. 11 D. 21

OpenStudy (anonymous):

-16+(-5_=-16-5=-21

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

is that d?

OpenStudy (anonymous):

or a?

OpenStudy (anonymous):

ups :S -16+(-5)=-16-5=-21, answer is A

OpenStudy (anonymous):

oh okay:)